OpenStudy (erniellyd28):

traffic signs are regulated by the manual on uniform traffic control devices (mutcd).the perimeter of a rectangular traffic sign is 72 inches.also its length is 6 inches longer than its width. find the dimensions of this sign.

jhonyy9 (jhonyy9):

rectangular so this mean that the perimeter will be equal 2l +2w than you know that the length is 6 inches longer than its width so what mean that l = w+6 in this case you can rewriting the P = 2l +2w = 2(w+6) +2w = 2w+12 +2w =4w+12 - but you know that P = 72 inches so than 72 = 4w+12 60=4w w=60/4 w=15 but you know that l=w+6 l=15+6 l=21 w=15 P=2l+2w 72=2*21 +2*15 72 =42+30 72=72 hope so much that this will help you sure right easy understanding good luck bye bye